OpenStudy (anonymous):

Tyree sold t-shirts for a Math Club fundraiser. During the first week, he sold 7 t-shirts, which was of his 35% inventory. How many t-shirts were in his inventory?

OpenStudy (anonymous):

Medal for best answer

OpenStudy (anonymous):

We don't know how many t-shirts were in his inventory, but we have to call it something, right? Let's call it x :) Tyree's sold 7 t-shirts, and this makes 35% of the all the t-shirts so 7 = 0.35x

OpenStudy (stamp):

I will not give you the answer, but we can solve the problem together.\[7=35\%i\] Can you solve for i?

OpenStudy (anonymous):

what do you mean both of you

OpenStudy (anonymous):

7 is 35% of all the shirts Tyree has. All Tyree has are x-shirts.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok carry on

OpenStudy (stamp):

\[tshirts=7\]\[tshirts=35\%\ inventory\]\[inventory=i\]\[35\%=\frac{35}{100}\]\[7=\frac{35}{100}i\]Mathematically, what is i equal to?

OpenStudy (anonymous):

the number of shirts

OpenStudy (anonymous):

I'm pretty sure stamp was asking for a numerical value :)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well since hes not here you get the medal but can you just give me the anser please?

OpenStudy (stamp):

@youarestupid absolutely not, this is a problem you are very well capable of solving. You have been presented setup, explanations, and support. If you cannot evaluate the terms me and @PeterPan have put forth for you, you may seriously be in the wrong class.

OpenStudy (anonymous):

I am in fourth grade and math is my bad subject.

OpenStudy (anonymous):

I am great in every thing else including computer programing.

OpenStudy (anonymous):

i am ok in spelling

OpenStudy (stamp):

That is fine. And I am glad to hear that you are interested in computer programming, it is a fine subject. It does involve a lot of mathematics though, many of the programs you create in college will involve far more advanced mathematical concepts. You have\[7=\frac{35}{100}x\]How do we change x from having a fraction to just being a single x? Remember, we are not interested in thirty-five hundredths of x, we are interested in what x is exactly equal to. Cross multiply? What do you obtain.

OpenStudy (anonymous):

You know a cool thing about equality? Whenever you have a = b You can multiply any number on both sides and still keep the = sign :D ka = kb Why don't we do that to \[7=\frac{35}{100}x\] ? Not a bad idea?

OpenStudy (anonymous):

oh i figurede it out it is 18

OpenStudy (anonymous):

sorry no it is 20

OpenStudy (stamp):

Ok so yes, you say twenty. Remember to write down\[20\ tshirts\]seeing as the question asked "how many tshirts". Now that we have our answer, even though we are right, it is good to know how to check. \[7=35\%(20)\]\[7=\frac{35}{100}20\]\[7=\frac{35(20)}{100}\]\[7=\frac{35}{5}\]\[7=7\] The math checks out, well done.

OpenStudy (anonymous):

bye